OpenStudy (geekfromthefutur):

The function below shows the number of students of a school who enrolled for cooking classes. Let f(x) represent the total number of students who enrolled for the classes after x years: f(x) = 11(1.35)x The average rate of change in the number of students who enrolled for cooking classes from the first to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (geekfromthefutur):

@satellite73 plz plz help me i really need your help i dont understand

OpenStudy (geekfromthefutur):

i give medal and fan

OpenStudy (anonymous):

is it \[f(x)=11(1.35)^x\]?

OpenStudy (geekfromthefutur):

yea :D

OpenStudy (anonymous):

i assume first year is actually year zero and fifth year is \(x=4\) your job is to compute \[\frac{f(4)-f(0)}{4-0}\]

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (anonymous):

i get this http://www.wolframalpha.com/input/?i=%2811*%281.35%29^4-11%29%2F4

OpenStudy (anonymous):

it is an average because the denominator is the number of years (4) and the numerator is the ending amount minus the initial amount

OpenStudy (geekfromthefutur):

oh ok but this isnt a multiple chose it says i have to round f(x)= 11(1.25)^x

OpenStudy (yanasidlinskiy):

What does that round up to, when you multiply?